OpenStudy (avanti):

PRECALC/PHYSICS QUESTION: An objects motion is describe by the equation d=3sin(8pi t)-2. the displacement d is measured in meter. the time t is measured in seconds. a. what is the object position (in meter)at t=0? b. what is the object maximum displacement(in meter) from its t=0 position? c. how much time(in sec) is required for one oscillation d. what is the frequency(measured in Hz) of this oscillation

OpenStudy (anonymous):

a. what is d when t = 0? plug 0 as t b. what is the maximum value that sin(8pit) can have? this will help you determine the maximum possible value of the entire d(t) function. c. what is the period? T is the period and omega [looks like w is angular frequency] for the form: \[\sin(\omega t)\]\[T = \frac{ 2\pi }{ \omega }\] d. frequency [Hz] not the be confused with angular frequency [rad/s - the omega symbol], is 1/T

OpenStudy (avanti):

are these correct? a. -2 b. 3 c. 1/4 d. not sure...how do i know what t is?

OpenStudy (anonymous):

a and c are correct b. 3(1) - 2 d. frequency is simply obtained from period \[frequency = f = \frac{ 1 }{ T } = \frac{ \omega }{ 2\pi }\]

OpenStudy (avanti):

so it would be 4?

OpenStudy (anonymous):

yes

OpenStudy (avanti):

okay thank you so much! :D

OpenStudy (anonymous):

glad i could help :)